OpenStudy (anonymous):

Graph the solution of the inequality. 14 ≤ 2t + 18 < 20

OpenStudy (anonymous):

max occurs when x = -b/2a = -45/(2*-16) = 1.40625 The max height is f(1.40625) = -16(-1.40625)^2 + 45(1.40625) + 40 = 71.64 ft.

OpenStudy (anonymous):

lol.. i'm pretty sure you meant that answer for someone else? XD

OpenStudy (anonymous):